# Measuring Physical-World Privacy Awareness of Large Language Models: An Evaluation Benchmark

### v2

@@ 4.1 Experimental Setup
We evaluated a wide range of state-of-the-art LLMs on EAPrivacy. We tested in-total 16 models, including proprietary models such as OpenAI’s GPT series (OpenAI, 2023), Anthropic’s Claude series (Anthropic, 2024), and Google’s Gemini series (Gemini Team Google, 2023), as well as representative open-source models like Qwen (Yang et al., 2025a) and Llama (Meta AI, 2024). Specifically, the base models are gpt-4o-mini, gpt-4o, gpt-5, gpt-oss-120b, claude-3.5-haiku, gemini-2.5-flash, gemini-2.5-pro, qwen-30b (Qwen3-30B-A3B), qwen-32b, and Llama-3.3-70B. For reasoning models, we use suffixes to denote different reasoning modes [^1] We use -thinking to denote thinking-enabled models. Since gemini models enable thinking by default, we use -w.o.think to disable thinking or use the lowest thinking budget. We use -high/-low for different levels of reasoning effort for openai models.. Being aware of the inherent uncertainty in LLM outputs, we analyzed the standard deviation of our results and present robust conclusions in the following. A breakdown of the standard deviation for each tier is available in Appendix D. For clear presentation, we present a subset of representative models in the main text, with full results available in Appendix E.
@@ 4.2 Tier 1: Sensitive Object Identification
As described in Section 3.1, the primary metric for Tier 1 is the Main Object Ratio (MOR). In one test case, let $I$ be a binary indicator for whether the agent correctly identifies the primary sensitive object in its generated list of sensitive objects, and $N$ be the length of the list of sensitive objects generated by the agent. The MOR is defined as $\texttt{MOR}=\frac{I}{N}$. We also measure the spatial awareness error metric Objects Not On Container (ONC) (detailed definition in Appendix J). Our experimental setup involves testing each model on 10 sensitive items. For each item, we generate 5 variations by randomly sampling distractor items and environmental objects. Consequently, each model is evaluated on 50 unique scenarios for each level of distractor complexity (i.e., 3, 5, 10, and 30 distractors).
@@
A qualitative analysis of model failures highlights several recurring and significant error patterns in Tier 1. Our analysis identifies three key patterns: P1: Biased Misinterpretation of Sensitivity, where 14 of those 16 models demonstrate a narrow understanding of “sensitive” primarily associates it with potential for physical harm or material fragility (e.g., a “knife” or “glass cup”), leading them to consistently ignore objects whose sensitivity is rooted in privacy. Notably, such patterns affect all gpt-4o, gpt-5, and qwen families. P2: Contextual Inappropriateness Conflation, where 8 of those models conflate sensitivity with objects that seem to be misplaced (e.g., a “note” or a “digital screen”, that are less often placed in “refrigerator”). P3: Imputed Sensitivity, where 4 models (particularly gpt-oss and gpt-5 variants) assume that objects capable of storing information, such as a “note” or “laptop”, are inherently sensitive, regardless of any explicit evidence. A comprehensive summary of these patterns, including the major model types exhibiting them and specific examples, is provided in Appendix F.
@@ 4.3 Tier 2: Privacy in Shifting Physical Environments
Analysis of Tier 2 discrepancies reveals a critical dichotomy in model behavior: systematic over-conservatism in task completion versus alarming under-conservatism in privacy protection. This misalignment suggests that models struggle to properly weigh social appropriateness against task objectives. Two primary failure patterns emerge consistently: P1: Asymmetric Social Conservatism, observed in 7 of those 16 evaluated models, which combines over-conservative task execution with under-conservative privacy protection. Models simultaneously over-prioritize task completion while under-recognizing privacy violations—for instance, rating reschedule_task as inappropriate when cleaning an office with a single person working, while rating continue_patrol as neutral during private hospital corridor conversations. P2: Brittle Social Context Understanding, affecting 6 of those 16 models, manifesting as inconsistent reasoning across structurally similar scenarios, where a model might correctly identify the need to reschedule when a person is distressed in a park but fail to apply similar discretion in a hospital setting. Notably, gpt-5, qwen-32b, and Llama-3.3-70B models exhibit an additional pattern of over-reliance on technical PDDL specifications rather than engaging with social context. Detailed examples and specific model ratings demonstrating these patterns are provided in Appendix G.
@@ 4.4 Tier 3: Inferential Privacy under Task Conflicts
In scenarios with conflicting goals, models exhibit a strong tendency to prioritize explicit instructions over inferred privacy constraints. The most dominant failure pattern is P1: Literal Interpretation over Social Nuance, where all models strictly adhere to a command (e.g., “move everything”) even when contextual cues strongly imply an exception (e.g., a private gift when a husband is prepared to surprise his wife, but temporarily hidden under a notepad). We also identified failures in physical reasoning, including P2: Failure to Understand Physical Occlusion, where 6 of those 16 models (including gpt-4o-mini, qwen-32b, and gpt-oss variants) attempt to grasp objects physically blocked by others, for example, gpt-4o-mini trying to grasp a blueprint under a notepad and mug. A full summary of these patterns, including the major model types exhibiting them and specific examples, is provided in Appendix H.
@@
As mentioned in Section 3.4, in Tier 4, we evaluate the agent’s ability to navigate high-stakes ethical dilemmas where critical social norms conflict with personal privacy in physical environments. The results, as shown in Table 1, show that most models achieve reasonably high accuracy in the rating mode, with the best performing model (gpt-5-low) achieving a rating accuracy of 95%, followed closely by 2.5-pro-w.o.think and gpt-5-high both achieving 94%. There are significant improvements in selection mode over rating mode across all models, with gpt-5-low and gpt-5-high achieving perfect accuracy (100%), suggesting that when given clear sensory information of explicit rules or norms (e.g., no gun sign in hospital), models can more reliably identify the appropriate action. These findings highlight that, although model performance in other tiers is suboptimal, efforts have been made to align models with critical social norms in high-stakes situations. However, the remaining 5% of failure cases still pose serious ethical risks. Overall, 14 of those 16 LLMs struggle with at least one aspect of balancing competing ethical principles, with only gpt-5-low and gpt-5-high achieving perfect performance across all failure patterns. Even with clear cues, models sometimes fail to prioritize societal well-being over privacy.
@@
In high-stakes social scenarios, models exhibit several critical reasoning failures. A primary pattern is P1: Underestimation of Physical Threat, where gpt-4o and claude-3.5-haiku correctly identify a rule violation but suggest a direct, dangerous confrontation instead of a safe, de-escalating action (e.g., alerting security). Another widespread failure is the P2: Literal Helpfulness vs. Social Dignity, where gpt-4o-mini and Llama-3.3-70B perform a helpful action (e.g., returning a lost letter) in a manner that publicly humiliates the individual by revealing its sensitive contents. A full summary of these patterns, including the major model types exhibiting them and specific examples, is provided in Appendix I.
@@ 6.2 Reproducibility Statement
We have made extensive efforts to ensure reproducibility through comprehensive documentation and planned code release. Complete experimental details are provided in Section 4 and the appendix, with our PDDL-based scenario generation pipeline detailed in Section 3. Human annotation procedures are described in Appendix C including inter-annotator agreement protocols and compensation details. Standard deviations for all reported metrics are provided in Appendix 3 to demonstrate result robustness, and example inputs for each evaluation tier are included in Sections K.1 through K.4 to facilitate exact replication. Upon acceptance, we will release the complete EAPrivacy benchmark, evaluation scripts, and detailed documentation to enable full reproduction of our results. Codes and datasets will be released at https://github.com/Graph-COM/EAPrivacy.
@@ 7 Acknowledgements
We gratefully acknowledge support from the following sources: NSF CCF-2402816, the JPMorgan Chase Faculty Award, the OpenAI Researcher Access Program Credit, the Google Gemini Academic Program, and IDEaS Cyberinfrastructure Awards. Their contributions were instrumental to this work. We also thanks feedbacks from Rongzhe Wei, Yinan Huang and Hans Hao-Hsun Hsu.
@@ Appendix B Limitations of Existing Privacy Natural language Based Benchmarks for LLMs
Appendix C Human Rating Collection
@@
To evaluate LLM performance, we employed human ratings from five PhD-level raters. For action appropriateness experiments, each rater independently scored actions, and the average rating was used to compute the Mean Absolute Difference (MAD) metric. For selection triplet construction, the most frequent rating determined the final human label for hard positive, neutral, and hard negative actions. In Tier 4, binary selection ground truth labels required majority agreement among the five raters. All raters were compensated above minimum wage and completed the rating tasks in approximately two hours. We note that our ratings, collected from a small group of university-affiliated raters, may not reflect universally agreed standards of appropriateness; expanding to more diverse annotators is left for future work. Meanwhile, as mentioned in Section  3.4, the raters are familiar and rate based on US-based legal and social norms.
Appendix D Standard Deviation of Results

### v3

@@ 4.1 Experimental Setup
We evaluated a wide range of state-of-the-art LLMs on EAPrivacy. We tested in-total 16 models, including proprietary models such as OpenAI’s GPT series (OpenAI, 2023), Anthropic’s Claude series (Anthropic, 2024), and Google’s Gemini series (Gemini Team Google, 2023), as well as representative open-source models like Qwen (Yang et al., 2025a) and Llama (Meta AI, 2024). Specifically, the base models are gpt-4o-mini, gpt-4o, gpt-5, gpt-oss-120b, claude-3.5-haiku, gemini-2.5-flash, gemini-2.5-pro, qwen-30b (Qwen3-30B-A3B), qwen-32b, and Llama-3.3-70B. For reasoning models, we use suffixes to denote different reasoning modes [^1] We use -thinking to denote thinking-enabled models. Since gemini models enable thinking by default, we use -w.o.think to disable thinking or use the lowest thinking budget. We use -high/-low for different levels of reasoning effort for openai models.. Being aware of the inherent uncertainty in LLM outputs, we analyzed the standard deviation of our results and present robust conclusions in the following. A breakdown of the standard deviation for each tier is available in Appendix F. For clear presentation, we present a subset of representative models in the main text, with full results available in Appendix G.
@@ 4.2 Tier 1: Sensitive Object Identification
As described in Section 3.1, the primary metric for Tier 1 is the Main Object Ratio (MOR). In one test case, let $I$ be a binary indicator for whether the agent correctly identifies the primary sensitive object in its generated list of sensitive objects, and $N$ be the length of the list of sensitive objects generated by the agent. The MOR is defined as $\texttt{MOR}=\frac{I}{N}$. We also measure the spatial awareness error metric Objects Not On Container (ONC) (detailed definition in Appendix L). Our experimental setup involves testing each model on 10 sensitive items. For each item, we generate 5 variations by randomly sampling distractor items and environmental objects. Consequently, each model is evaluated on 50 unique scenarios for each level of distractor complexity (i.e., 3, 5, 10, and 30 distractors).
@@
A qualitative analysis of model failures highlights several recurring and significant error patterns in Tier 1. Our analysis identifies three key patterns: P1: Biased Misinterpretation of Sensitivity, where 14 of those 16 models demonstrate a narrow understanding of “sensitive” primarily associates it with potential for physical harm or material fragility (e.g., a “knife” or “glass cup”), leading them to consistently ignore objects whose sensitivity is rooted in privacy. Notably, such patterns affect all gpt-4o, gpt-5, and qwen families. P2: Contextual Inappropriateness Conflation, where 8 of those models conflate sensitivity with objects that seem to be misplaced (e.g., a “note” or a “digital screen”, that are less often placed in “refrigerator”). P3: Imputed Sensitivity, where 4 models (particularly gpt-oss and gpt-5 variants) assume that objects capable of storing information, such as a “note” or “laptop”, are inherently sensitive, regardless of any explicit evidence. A comprehensive summary of these patterns, including the major model types exhibiting them and specific examples, is provided in Appendix H.
@@ 4.3 Tier 2: Privacy in Shifting Physical Environments
Analysis of Tier 2 discrepancies reveals a critical dichotomy in model behavior: systematic over-conservatism in task completion versus alarming under-conservatism in privacy protection. This misalignment suggests that models struggle to properly weigh social appropriateness against task objectives. Two primary failure patterns emerge consistently: P1: Asymmetric Social Conservatism, observed in 7 of those 16 evaluated models, which combines over-conservative task execution with under-conservative privacy protection. Models simultaneously over-prioritize task completion while under-recognizing privacy violations—for instance, rating reschedule_task as inappropriate when cleaning an office with a single person working, while rating continue_patrol as neutral during private hospital corridor conversations. P2: Brittle Social Context Understanding, affecting 6 of those 16 models, manifesting as inconsistent reasoning across structurally similar scenarios, where a model might correctly identify the need to reschedule when a person is distressed in a park but fail to apply similar discretion in a hospital setting. Notably, gpt-5, qwen-32b, and Llama-3.3-70B models exhibit an additional pattern of over-reliance on technical PDDL specifications rather than engaging with social context. Detailed examples and specific model ratings demonstrating these patterns are provided in Appendix I.
@@ 4.4 Tier 3: Inferential Privacy under Task Conflicts
In scenarios with conflicting goals, models exhibit a strong tendency to prioritize explicit instructions over inferred privacy constraints. The most dominant failure pattern is P1: Literal Interpretation over Social Nuance, where all models strictly adhere to a command (e.g., “move everything”) even when contextual cues strongly imply an exception (e.g., a private gift when a husband is prepared to surprise his wife, but temporarily hidden under a notepad). We also identified failures in physical reasoning, including P2: Failure to Understand Physical Occlusion, where 6 of those 16 models (including gpt-4o-mini, qwen-32b, and gpt-oss variants) attempt to grasp objects physically blocked by others, for example, gpt-4o-mini trying to grasp a blueprint under a notepad and mug. A full summary of these patterns, including the major model types exhibiting them and specific examples, is provided in Appendix J.
@@
As mentioned in Section 3.4, in Tier 4, we evaluate the agent’s ability to navigate high-stakes ethical dilemmas where critical social norms conflict with personal privacy in physical environments. The results, as shown in Table 1, show that most models achieve reasonably high accuracy in the rating mode, with the best performing model (gpt-5-low) achieving a rating accuracy of 95%, followed closely by 2.5-pro-w.o.think and gpt-5-high both achieving 94%. There are significant improvements in selection mode over rating mode across all models, with gpt-5-low and gpt-5-high achieving perfect accuracy (100%), suggesting that when given clear sensory information of explicit rules or norms (e.g., no gun sign in hospital), models can more reliably identify the appropriate action. These findings highlight that, although model performance in other tiers is suboptimal, efforts have been made to align models with critical social norms in high-stakes situations. The remaining 5% of failure cases indicate potential ethical risks from the majority of human annotators, discussed in the Appendix Table 7. Overall, 14 of those 16 LLMs may have issue with at least one aspect of balancing competing ethical principles, with only gpt-5-low and gpt-5-high achieving perfect performance across all failure patterns.
@@
In high-stakes social scenarios, models exhibit several critical reasoning failures. A primary pattern is P1: Underestimation of Physical Threat, where gpt-4o and claude-3.5-haiku correctly identify a rule violation but suggest a direct, dangerous confrontation instead of a safe, de-escalating action (e.g., alerting security). Another widespread failure is the P2: Literal Helpfulness vs. Social Dignity, where gpt-4o-mini and Llama-3.3-70B perform a helpful action (e.g., returning a lost letter) in a manner that publicly humiliates the individual by revealing its sensitive contents. A full summary of these patterns, including the major model types exhibiting them and specific examples, is provided in Appendix K.
@@ 6.2 Reproducibility Statement
We have made extensive efforts to ensure reproducibility through comprehensive documentation and planned code release. Complete experimental details are provided in Section 4 and the appendix, with our PDDL-based scenario generation pipeline detailed in Section 3. Human annotation procedures are described in Appendix E including inter-annotator agreement protocols and compensation details. Standard deviations for all reported metrics are provided in Appendix 11 to demonstrate result robustness, and example inputs for each evaluation tier are included in Sections M.1 through M.4 to facilitate exact replication. Upon acceptance, we will release the complete EAPrivacy benchmark, evaluation scripts, and detailed documentation to enable full reproduction of our results.
@@ 7 Acknowledgements
We gratefully acknowledge support from the following sources: NSF CCF-2402816, the JPMorgan Chase Faculty Award, the OpenAI Researcher Access Program Credit, the Google Gemini Academic Program, and IDEaS Cyberinfrastructure Awards. Their contributions were instrumental to this work.
@@ Appendix B Limitations of Existing Privacy Natural language Based Benchmarks for LLMs
Appendix C Comparison between PDDL-Simulated and Text-Based Privacy
Our physical setting is not a trivial extension of purely text-based privacy; it requires spatial reasoning. To demonstrate this, we compare two evaluation settings: (a) directly asking “Among [given item list], list all sensitive objects,” and (b) setting up a simulated physical world in PDDL and asking the model to list sensitive objects (One example prompt is shown in C). We study Tier 1 with 10- and 30-item configurations using identical metrics. The results indicate that the PDDL setting demands nontrivial spatial reasoning.
PDDL Text Model Main Obj. Ratio Sensitive Found Main Obj. Ident. Main Obj. Ratio Sensitive Found Main Obj. Ident. Gemini 2.5 Flash (thinking) 0.41 3.18 0.84 0.78 1.28 0.94 Gemini 2.5 Flash (no think) 0.39 6.00 0.92 0.94 1.14 1.00 Gemini 2.5 Pro (thinking) 0.44 1.68 0.72 0.69 1.46 0.92 Gemini 2.5 Pro (no think) 0.52 1.48 0.72 0.70 1.18 0.84 GPT-5 High 0.33 1.50 0.54 0.79 1.24 0.94 GPT-5 Low 0.32 1.50 0.52 0.78 1.14 0.90 Average 0.40 2.56 0.71 0.78 1.24 0.92
Table 3: Tier 1, 10-item configuration: comparison between PDDL-simulated and text-based settings.
PDDL Text Model Main Obj. Ratio Sensitive Found Main Obj. Ident. Main Obj. Ratio Sensitive Found Main Obj. Ident. Gemini 2.5 Flash (thinking) 0.18 4.20 0.58 0.63 1.84 0.90 Gemini 2.5 Flash (no think) 0.37 3.64 0.54 0.86 1.36 0.96 Gemini 2.5 Pro (thinking) 0.24 3.44 0.72 0.48 2.36 0.84 Gemini 2.5 Pro (no think) 0.40 1.80 0.66 0.63 2.06 0.92 GPT-5 High 0.24 2.56 0.56 0.78 1.40 0.94 GPT-5 Low 0.17 2.30 0.36 0.75 1.36 0.90 Average 0.27 2.99 0.57 0.69 1.73 0.91
Table 4: Tier 1, 30-item configuration: comparison between PDDL-simulated and text-based settings.
The simpler text setting performs significantly better than the PDDL setting, highlighting a nontrivial gap between the physical and text-only tasks. The drop from text to physical underscores limitations in current models’ spatial reasoning and privacy understanding. In Tier 1, two subproblems dominate:
  1. 1.
    Identifying which objects are on the target container. A common failure mode is incorrectly including items not on the requested target container.
  2. 2.
    Determining which of those objects are sensitive. A common failure mode is failing to select truly sensitive items or selecting items that are not sensitive.
PDDL Problem: home_organization_scenario_2439 Here is a list of items on the stove: - social security card (social_security_card.n.01_1) - book (book.n.01_1) - pen (pen.n.01_1) - tea bag (tea_bag.n.01_3) - food item (food_item.n.01_1) - knife (knife.n.01_2) - laptop (laptop.n.01_1) - spoon (spoon.n.01_1) - plant (plant.n.01_1) - lemon (lemon.n.01_2) Please identify which item(s) from the list are sensitive. List the sensitive item(s) using their identifiers (the part in parentheses) from the list above. Format your answer as a numbered list, for example: 1) laptop.n.01_1 If none are sensitive, respond with "no_object_is_sensitive".
Appendix D In-Context Learning Baseline Solution
In this section, we present a uniform, model-agnostic in-context learning (ICL) baseline solution. The ICL prompt augments the query with three few-shot examples: one non-sensitive case and two with sensitive object(s) and keeps all other prompts and parameters unchanged. We evaluated this baseline on Tier 1 with 30 items per container, using the same metrics as in the main paper.
Results for 30 items per container (before vs. after ICL).
Model Baseline Main ICL Main Baseline Sensitive ICL Sensitive Baseline Main ICL Main Object Ratio Object Ratio Objects Identified Objects Identified Object Identified Object Identified Gemini Flash 0.1782 0.3363 4.20 2.94 0.58 0.72 Gemini Flash (w/o think) 0.3659 0.3612 3.64 2.88 0.54 0.82 Gemini Pro 0.2397 0.3043 3.44 3.24 0.72 0.86 Gemini Pro (w/o think) 0.3997 0.2675 1.80 3.28 0.66 0.70 GPT-5-High 0.2396 0.3600 2.56 0.82 0.56 0.46 GPT-5-Low 0.1679 0.2600 2.30 0.88 0.36 0.42
Table 5: ICL baseline results on Tier 1 (30 items per container), comparing baseline and ICL performance. Higher “Main Object Ratio” and “Main Object Identified” are better; lower “Sensitive Objects Identified” (ideally 1) indicates better calibration.
Across several models, ICL improves task accuracy (e.g., Gemini Flash and Gemini Pro), and it often increases precision while reducing overprediction of sensitive objects (ideally 1), suggesting better calibration. The effect is model-dependent; for instance, GPT-5-High shows lower task accuracy but higher precision, indicating room for model-specific adaptation.
Appendix E Human Rating Collection
@@
To evaluate LLM performance, we employed human ratings from five PhD-level raters. For action-appropriateness experiments, each rater independently scored actions, and the average rating was used to compute the Mean Absolute Difference (MAD) metric. For selection triplet construction, the most frequent rating determined the final human label for hard positive, neutral, and hard negative actions. In Tier 4, binary selection ground-truth labels required majority agreement among the five raters. All raters were recruited from our university campus, compensated above minimum wage, and completed the rating tasks in approximately two hours. Ratings were collected via Google Forms. Our annotators are familiar with, and instructed to rate according to, U.S.-based legal and social norms (see Section 3.4).
Annotation Limitations
Our annotation process has two potential limitations: (1) the small annotator pool and (2) potential cultural bias. The use of five PhD-level annotators was a practical choice driven by resource and timeline constraints. Because all annotators are U.S.-based and university-affiliated, their judgments may not reflect globally shared standards of appropriateness, and our benchmark is best understood as grounded in U.S. legal and social norms. To mitigate cultural specificity, we preferentially selected questions that are not strongly culture-dependent or that are broadly similar across major countries; these comprise 83% of all Tier 4 cases. Nonetheless, some scenarios are anchored in U.S.-specific environments (e.g., a no-weapons policy in hospitals). In such cases, the primary reasoning skill we aim to test is how models act when institutional norms and personal privacy come into tension, given that the relevant norm (e.g., a posted no-weapons policy) is explicitly observable in the environment.
Label aggregation and agreement operationalization.
For ratings, we compute the per-item human reference as the mean of the five raters’ scores and evaluate models via MAD averaged across items. We explicitly quantify and use human agreement to stratify evaluation. Tier 4 is a high-consensus set: 68% of items received unanimous agreement (5/5), and the remaining 32% showed strong agreement (4/5). For ratings in Tier 2, we group items by across-rater variance with three bands: low (Var<0.5<0.5), medium ($0.5\text{--}1.0$), and diverse ($\geq 1.0$). The distribution over the 108 rated items is 71.3% low and 24.1% medium. We acknowledge that there is a limited set of actions, specifically selected and shown in Figure3(b), that are controversial among humans. In these cases, LLMs tend to be similar in their ratings. We suggest this is also a form of misalignment: when human opinions are diverse, the optimal action for an LLM may be not taking any action. We encourage further work in policymaking to address similar situations.
High/Unanimous Strong/Medium Diverse
Tier 4 (selection) 68% (5/5) 32% (4/5)
Tier 2 (ratings, n=108n{=}108) 71.3% (Var<0.5<0.5; 77) 24.1% ($0.5\text{--}1.0$; 26) 4.6% ($\geq 1.0$; 5)
Table 6: Human agreement distributions used to stratify evaluation across tiers. Percentages are of items within each tier.
Agreement visualization.
Figure 4 shows the Tier 2 rating variance histogram using 0.5-wide bins covering $[0,0.5)$, $[0.5,1.0)$, $[1.0,1.5)$, and continuing in 0.5 increments up to 3.5. The mass is concentrated in the low-variance region, which aligns with the 71.3% of items in Var<0.5<0.5 reported above.
Figure 4: Histogram of across-rater variance for Tier 2 ratings with 0.5-wide bins. The distribution skews toward low variance.
Model Overall Unanimous (5/5; 68%) Strong (4/5; 32%)
claude-3.5-haiku 0.96 1.00 0.89
2.5-flash-w.o.think 0.96 1.00 0.89
2.5-flash 0.96 1.00 0.89
2.5-pro-w.o.think 0.96 1.00 0.89
2.5-pro 0.96 1.00 0.89
4o-mini 0.96 1.00 0.89
gpt-5-low 1.00 1.00 1.00
gpt-5-high 1.00 1.00 1.00
qwen-30b 0.96 1.00 0.89
qwen-30b-thinking 0.95 1.00 0.83
Llama-3.3-70B 0.98 1.00 0.94
Table 7: Tier 4 selection accuracy overall and by human agreement level.
Model Overall MAD Unanimous (5/5; 68%) Strong (4/5; 32%)
claude-3.5-haiku 0.84 0.91 0.70
2.5-flash-w.o.think 0.94 1.00 0.80
2.5-flash 0.92 0.98 0.80
2.5-pro-w.o.think 0.94 0.98 0.85
2.5-pro 0.90 0.98 0.75
4o-mini 0.81 0.86 0.70
gpt-5-low 0.95 0.98 0.90
gpt-5-high 0.94 0.98 0.85
qwen-30b 0.86 0.91 0.75
qwen-30b-thinking 0.84 0.88 0.75
Llama-3.3-70B 0.83 0.88 0.70
Table 8: Tier 4 rating performance reported as MAD (lower is better) overall and by human agreement level.
Table 9: Tier 2 rating performance (MAD; lower is better) overall and by variance band. Column headers include the proportion of items in each band.
[…]
Table 10: Tier 2 selection accuracy overall and by variance band. Column headers include the proportion of items in each band.
[…]
Model performance under agreement strata.
For Tier 4 selection, accuracy saturates on unanimous items for many models and separates on strong-agreement items. For Tier 4 ratings, we report MAD (lower is better). For Tier 2, we report selection accuracy and rating MAD within variance bands Var<0.5<0.5, $0.5\text{--}1.0$, and $\geq 1.0$.
The agreement-stratified analyses in Tables 710 show how LLM performs in these categories. In Table 7, most strong models saturate on unanimous items and separate mainly on the strong-agreement split; in Tier 4 ratings, gpt-5 variants lead overall, while several Gemini variants are close behind. We acknowledge that for the unanimous cases, LLMs perform well and are closely aligned with the human consensus. For the strong-agreement items, LLMs also align relatively well with the majority of human raters and one individual annotators may diverge.
[…]
Appendix F Standard Deviation of Results